# paper1.bib

### v2

@@ 1 Introduction
The aim of this paper is to investigate the conditions under which a quasi-Newton method can achieve finite termination on (QP) without the need for exact line search. In Section 2 we briefly discuss Krylov subspaces and Krylov subspace methods such as the method of conjugate gradients as well as quasi-Newton methods. In Section 3 we introduce the concept of a subspace Newton step within a given Krylov subspace as well as an expanding Krylov subspace. We generalize to a subspace quasi-Newton step in Section 4, showing that it suffices to create a quasi-Newton matrix based on at most two vectors. Finally we present an iterative memoryless quasi-Newton method based on at most two vectors, in addition to the gradient, that will compute the Newton step in a finite number of iterations with the use of arbitrary step sizes in Section 5.
@@ 5 An exact quasi-Newton method without exact line search
Theoretically, if we have explicit knowledge of $H$ then even a zero step will lead to generation of the Newton step as iteratively new conjugate directions are calculated and then Newton scaled with respect to the fixed point. Without explicit knowledge of $H$ it is vital to take a nonzero step in order to evaluate a new gradient. It is then possible to use a difference of gradients to learn how $H$ acts on the step taken. This is described in Lemma A.1.
@@ 6 Conclusion
Appendix A Appendices
A.1 A first order formulation of Algorithm 1
The construction of $B_{k+1}$ in Algorithm 1 requires information on how $H$ acts on the vectors $q_{k}$ and $p^{N}_{k}$. When a nonzero step is taken at each iteration, this information can be gathered using a difference of gradients. Explicit knowledge of the Hessian $H$ is unnecessary.
Lemma A.1
$Hq_{k}=\frac{1}{\alpha_{k}}(g_{k+1}-g_{k})-Hp^{N}_{k-1}$ (A.1)
$p^{N}_{k}=\left(-\frac{\alpha_{k}g_{k}^{T}q_{k}}{q_{k}^{T}(g_{k+1}-g_{k})}-1\right)q_{k}+(1-\alpha_{k})p_{k}$ (A.2)
$Hp^{N}_{k}=\left(-\frac{\alpha_{k}g_{k}^{T}q_{k}}{q_{k}^{T}(g_{k+1}-g_{k})}-1\right)Hq_{k}+\left(\frac{1}{\alpha_{k}}-1\right)(g_{k+1}-g_{k}).$ (A.3)
When $k\geq r$, $p^{N}_{k}=(1-\alpha_{k})p_{k}$, so
$Hp^{N}_{k}=\left(\frac{1}{\alpha_{k}}-1\right)(g_{k+1}-g_{k}).$ (A.4)
$\frac{1}{\alpha_{k}}(g_{k+1}-g_{k})=Hp_{k}=Hq_{k}+Hp^{N}_{k-1}.$
Given that the matrix $B_{k}$ acts as $H$ on $p^{N}_{k-1}$, the formulation of $Hq_{k}$ in (A.1) follows directly.
$=\frac{1}{\alpha_{k}}q_{k}^{T}(g_{k+1}-g_{k})$
in (5.1) to obtain the expression (A.2). The formulations for $Hp^{N}_{k}$ in (A.3) and (A.4) follow immediately.
Using the expressions for $p^{N}_{k},Hq_{k}$ and $Hp^{N}_{k}$ given by the lemma above means that Algorithm 1 can be implemented using first order information only, with the same termination guarantees.
A.2 Some linear algebra results

### v3

@@ 1 Introduction
The aim of this paper is to investigate the conditions under which a quasi-Newton method can achieve finite termination on (QP) without the need for exact line search. In Section 2 we briefly discuss Krylov subspaces and Krylov subspace methods such as the method of conjugate gradients as well as quasi-Newton methods. In Section 3 we introduce the concept of a subspace Newton step within a given Krylov subspace as well as an expanding Krylov subspace. We generalize to a subspace quasi-Newton step in Section 4, showing that it suffices to create a quasi-Newton matrix based on at most two vectors. Finally we present an iterative memoryless quasi-Newton method based on at most two vectors, in addition to the gradient, that will compute the Newton step in a finite number of iterations with the use of arbitrary step sizes in Section 5. Additionally, in Section 5, numerical support for the termination guarantees and a first order only formulation of the algorithm are provided.
@@ 5 An exact quasi-Newton method without exact line search
Theoretically, if we have explicit knowledge of $H$ then even a zero step will lead to generation of the Newton step as iteratively new conjugate directions are calculated and then Newton scaled with respect to the fixed point. Without explicit knowledge of $H$ it is vital to take a nonzero step in order to evaluate a new gradient. It is then possible to use a difference of gradients to learn how $H$ acts on the step taken. This is described in Subsection 5.2.
@@
5.1 Numerical support
Given a fixed initial point $x_{0}$, it is known that BFGS and CG terminate in exactly $r$ iterations when exact line search is used. Conversely, Algorithm 1 will terminate in at most $r+1$ iterations as long as a unit step size is chosen at the last iteration by Corollary 5.1. The constant $r$ is the dimension of the largest Krylov subspace $\mathcal{K}_{r}(g_{0},H)$, or equivalently the number of distinct eigenvalues of $H$ such that the initial gradient $g_{0}$ is not orthogonal to the corresponding eigenspace.
For numerical support, tests were implemented with varying values of $r$. To mimic the results given using exact arithmetic, the implementation was run in Matlab using high precision of $64$ digits and low dimensions $n=20,40$. For termination, a tolerance criterion for the norm of the gradient was set to sqrt(eps). The Hessian $H$ and vector $c$ were constructed as
$H=\mathop{\operator@font diag}\left(\!\begin{array}[]{cccccc}1,&\ldots,&r,&1,&\ldots&n-r\end{array}\!\right),\qquad c=\left(\!\begin{array}[]{cccc}1,&1,&\ldots,&1\end{array}\!\right)^{T},$
where $r\geq n/2$. With the initial point $x_{0}$ at the origin, there are exactly $r$ distinct eigenvalues of $H$ with eigenspaces not orthogonal to the gradient $g_{0}$. This guarantees that in theory BFGS and CG will terminate at iteration $r$, and Algorithm 1 terminates in $r+1$ iterations when $\alpha_{r}=1$ by Corollary 5.1. For all iterations of the implementation of Algorithm 1, the constant $\sigma_{k}$ was set to $1$.
Termination
Algorithm 1
$n$ $r$ BFGS CG $\alpha_{k}=0$ $\alpha_{k}=1$ $\alpha_{k}\sim U(0,1)$
20 10 10 10 11 11 11
15 15 15 16 16 16
20 20 20 21 21 21
40 20 20 20 21 21 21
30 30 30 31 31 31
40 40 40 41 41 41
Table 1: Termination of algorithms on quadratic problems
The termination results are shown in Table 1. The step sizes given for Algorithm 1 in Table 1 reference the choice of $\alpha_{k}$ for $k<r$. The algorithm was implemented using constant step sizes $\alpha_{k}=0,1$ as well as $\alpha_{k}$ sampled from the uniform distribution $U(0,1)$. For $k\geq r$, unit steps are taken until termination. The two methods BFGS and CG terminate in the expected $r$ iterations with exact line search. For all choices of step sizes for Algorithm 1, termination occurs in exactly $r+1$ iterations as expected from Corollary 5.1.
5.2 A first order formulation of Algorithm 1
The construction of $B_{k+1}$ in Algorithm 1 requires information on how the Hessian $H$ acts on the vectors $q_{k}$ and $p^{N}_{k}$. It is however unnecessary to have explicit knowledge of $H$. While the formulation of Algorithm 1 allows for any step sizes $\alpha_{k}$, to use only first order information requires that new gradients be evaluated at each iteration. Therefore it is necessary to choose $\alpha_{k}\neq 0$. When a nonzero step is taken at each iteration, this information can be gathered using a difference of gradients.
Lemma 5.1
$Hq_{k}=\frac{1}{\alpha_{k}}(g_{k+1}-g_{k})-Hp^{N}_{k-1}$ (5.2)
$p^{N}_{k}=\left(-\frac{\alpha_{k}g_{k}^{T}q_{k}}{q_{k}^{T}(g_{k+1}-g_{k})}-1\right)q_{k}+(1-\alpha_{k})p_{k}$ (5.3)
$Hp^{N}_{k}=\left(-\frac{\alpha_{k}g_{k}^{T}q_{k}}{q_{k}^{T}(g_{k+1}-g_{k})}-1\right)Hq_{k}+\left(\frac{1}{\alpha_{k}}-1\right)(g_{k+1}-g_{k}).$ (5.4)
When $k\geq r$, $p^{N}_{k}=(1-\alpha_{k})p_{k}$, so
$Hp^{N}_{k}=\left(\frac{1}{\alpha_{k}}-1\right)(g_{k+1}-g_{k}).$ (5.5)
$\frac{1}{\alpha_{k}}(g_{k+1}-g_{k})=Hp_{k}=Hq_{k}+Hp^{N}_{k-1}.$
Given that the matrix $B_{k}$ acts as $H$ on $p^{N}_{k-1}$, the formulation of $Hq_{k}$ in (5.2) follows directly.
$=\frac{1}{\alpha_{k}}q_{k}^{T}(g_{k+1}-g_{k})$
in (5.1) to obtain the expression (5.3). The formulations for $Hp^{N}_{k}$ in (5.4) and (5.5) follow immediately.
Using the expressions for $p^{N}_{k},Hq_{k}$ and $Hp^{N}_{k}$ presented in the lemma above means that Algorithm 1 can be implemented using first order information only to recover the action of $H$ on the necessary vectors. The termination guarantees of the algorithm remain unchanged with this formulation.
@@ 6 Conclusion
Appendix A Appendix: Some linear algebra results